OpenStudy (anonymous):

make x the subject of the following x+3=2a+bx

OpenStudy (anonymous):

x+3=2a+bx First group the x on one side x-bx=2a-3 Now factor the left x(1-b)=2a-3 divide x=(2a-3)/(1-b)

OpenStudy (anonymous):

x-bx=2a-3===> x(1-b)=2a-3 ===> x= )2a-3)/(1-b)